Operations Report – January 29th

Matt Wise
XO/Engineer
MDRS Crew 220

Operations Report

Crew 220 Operations Report 29JAN2020

Sol: 10

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater, Hab lower deck 1 fluorescent light ballast.

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 13hrs
From what time last night: 1800

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 81%

Diesel Reading – 1/3

Station Propane Reading- 80%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0148498,4

Water (static tank) (gal) – 123

Static to loft Pump used – Yes 38 gal

Water in GreenHab (gal): 127.94

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Deimos rover used: Still in the workshop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Yes

Starting Hours: 128.3

Beginning Charge: 100%

Ending Hours: 128.6

Ending Charge: 63%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 134.3

Beginning Charge: 100%

Ending Hours: 134.6

Ending Charge: 62%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? yes, water and mail

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: 3 radios malfunctioning
Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck malfunctioning. The problem seems to stem from irregular current flow from temp generator. Center light bar has been fixed and seems to function properly at all times, light bar near back airlock has been repaired and seems to be functioning properly, while light bar near stairs has no bulbs.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

Questions, concerns, and requests to Mission Support: n/a

Greenhab Report – January 29th

  


 

GreenHab Officers: Marufa Bhuiyan and Morgan Kainu

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on.

Average temperatures: 22.7°C

09:56

Floor Unit: 22 °C

12:58

Floor Unit: 25 °C

19:09

Floor Unit: 21 °C

GreenHab door opened: 12:58 and closed at 13:30.

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 8.7 gallons

Daily water usage for research and/or other purposes: NA

Water in Blue Tank 127.94 gallons

Time(s) of watering for crops:

10:04 = 6.6 Gal

19:10 = 2.1 Gal

Changes to crops: Re-arranged some plants in the GreenHab according to their needs (e.g. sunlight, heat, watering schedule etc.)

Planted: 1 g radish sparkler, 0.3 g onion evergreen white bunching

Harvest: 77 g icicle radish, 10 g purple sun carrots, 6 g lettuce, 1 g chives, 1 g cilantro, 3 g rosemary, 2 g butter crunch, 2 g oregano, 11 g butter brunch, 30 g green leaf chard, 4 g lettuce

Support/supplies needed: Small disposal until or, recycle bin/trash basket

Other Notes:

Added fertilizer to some plants, and a new seed bank added to GreenHab.

Ginger is being prepared for planting in a dish; waiting for sprouts before planting.

Changed the batteries in the scale.

Added hand sanitizer and paper towels to GreenHab.

EVA Report – January 29th

  

Crew 220 (MAU) EVA Report 29JAN2020

EVA 10

Author: Johannes Svensoy, MD

Purpose of EVA: Exploratory and scouting EVA

Start time: 13:30

End time: 16:30

Narrative: The crew members going for the EVA started at MAU station and walked up Cow Dung Road over to MDRS Station. As previously arranged, the crew took the rovers Spirit and Curiosity for the exploratory EVA going North. The crew drove North on Cow Dung Road to Cowboy Corner. From there, the crew members walked East by foot and were able to ascend the hills to appreciate sprawling vistas. We were able to scout further North and take down coordinates important to help map the area. Everything went just as planned and the crew came back in good health.

Destination: Cowboy Corner.

Coordinates (use UTM NAD27 CONUS): 4252500, 518500

Participants: HabComm: Connie (MDRS). EVA Crew: Shawna, Morgan, Lee, and Johannes.

Road(s) and routes per MDRS Map: Cow Dung Road 0110 North passing North Ridge to Cowboy Corner.

Mode of travel: Rovers Spirit and Curiosity and by foot.

Support Requested: No support requested during EVA.

Sol Summary – January 29th

Crew 220 (MAU) Sol 10 Summary 29Jan2020

Summary Title: All Aboard the Sol 10!
Author’s name: Shawna Pandya, Commander
Mission Status: MAUtastic
Sol Activity Summary: We had an extremely productive day at MAU. The morning consisted of in-depth upkeep, cleaning, and maintenance, and we achieved a lot in getting our electrical operations, data network and station cleanliness up to par. In other news, we hosted our first-ever visiting Ambassador from MDRS, Officer Kainu, and deem our first exchange extremely successful, and look forward to many more. The afternoon saw a successful exploratory EVA to Cowboy’s Corner with the MAU Crew, followed by an evening sojourn to MDRS for a cultural exchange including martial arts and dance demonstrations, and culminating with a "Future Cafe."
Look Ahead Plan: As our Mission winds down, we will be filling the remaining Sols with medical technology demonstrations, academic activities, wellness and a few more EVAs.
Anomalies in work: None
Weather: -6C low / 5C max, 0C average.
Crew Physical Status: MAUminal
EVA: Successful Exploratory EVA to Cowboy’s Corner completed.
Reports to be filed:
Sol Summary
Commander’s Report
Photo Report
Operations Report
EVA Report
EVA Request: Pending.
Support Requested: If you could invent a mud-cleaning machine, MAU Station would LOVE to hear about it! 🙂

Commander Report – January 29th

Crew 220 (MAU) Commander’s Report 29Jan2020

Author: Shawna Pandya, Commander

Title: Some Sol-Searching…

We had an extremely productive day at MAU today. We hosted our first-ever MDRS Ambassador, MDRS Science Officer Kainu, overnight at MAU as our first-ever guest, and had a delightful exchange of ideas and competencies. All in all, I would call this first exchange a success, and am keen to continue to host ambassadors during the rest of the mission to continue the exchange of skills, competencies, and cultural aspects of the two stations.

On the operations and upkeep side, the MDRS and MAU Crew Engineers whipped up some Mars-gic and reinforced our power grid with our online and back-up power systems. The MAU Engineer reinforced the structural integrity of the station, while the MDRS Engineer worked on establishing a wireless station-to-station data link, and will continue to do so over the next ensuing Sols. The MAU Crew as a whole came together to reinforce the station flooring, reorganize the station, and scrub it down. It was hard work, but incredibly rewarding to see our station shine from our labours!

The day ended with a MAU exploratory EVA to Cowboy Corner, a special little part of Mars that affords views of dazzling expanses of marbled reds, whites and greys that rise to meet the eye to show off their grandeur in full force. Not many humans are privy to such a grand yet secret spectacle, and we are indeed privileged to count ourselves amongst them.

The Sols, they wind down now, and we will make the most of every last Martian minute.

Until my next communication,

Commander Pandya
Callsign: Nightowl
MAU Station

Operations Report – January 29th

Crew 220 (MAU) Operations Report 29JAN2020

SOL 10

Name of person filing report: Lee Roberts, Crew Engineer

Non-nominal systems: N/A

Generator: Operating nominally.

Hours run: Not run overnight

From what time last night: N/A

To what time this morning: N/A

List of any additional daytime hours when the generator was run: 2 hours

Solar- SOC% (Before the generator is run at night): N/A

Diesel Reading – N/A

Station Propane Reading- 100% + Backup (1.8 cylinders)

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – N/A

Water Meter (units)- N/A

Water (static tank) (gal) – N/A

Static to loft Pump used – N/A

Water in GreenHab (gal): N/A

Water in Science Dome (gal): N/A

Toilet Tank Emptied – no

Deimos rover used: N/A

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Used for exploration EVA

Starting Hours: 128.3

Beginning Charge: 100%

Ending Hours: 128.6

Ending Charge: 63%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging: Yes

Curiosity Rover used: Used for exploration EVA

Starting Hours: 134.3

Beginning Charge: 100%

Ending Hours: 134.6

Ending Charge: 62%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used to town to get water.

CrewCar used and why, where? Not used.

General Notes and Comments:

Summary of the internet: N/A

Summary of Suit and Radios: Nominal

Summary of Hab: The crew spent several hours this morning cleaning the Station and trying to mitigate mud intrusion on the floors. Tarps were placed on the floor in two of the rooms to reinforce the floor surfaces.

During external inspection of the Station it was found that several supporting poles for the Station walls were dislodged during the recent windstorm, compromising the structural integrity. These poles were reaffixed, and the walls appear to be structurally sound.

In addition to replenishing the propane supply, we also installed a charger to maintain our backup battery at a steady charge, and we are instituting a power management plan to reduce propane consumption on the generator.

Summary of Science Dome operations: Temperatures are hovering around 5°C during the day.

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None

Journalist Report – January 29th

Crew 220 MDRS Journalist Report 29JAN2020
Author: Morgan Kainu, Crew Journalist
Sol 10
Title: Destination Exploration

Body: Today marked the first day of the MAU Ambassador Program implementation. A crew member from MDRS was hosted at the MAU hab the previous night to learn from the MAU crew about internal operations as well as to stay the day and to participate with the MAU crew in an explorative extravehicular activity (EVA). This was the first EVA to leave directly from the MAU hab.
It was a mildly warm day, yet stunning red Martian day. It was perfect for exploring and admiring the landscape. The MAU crew and MDRS ambassador crew member together expressed their sentiments about the extreme beauty experienced on the EVA, reflecting in their mind’s eye. They also took this opportunity to share their photos from the day with the rest of the MDRS crew back at their hab.

Both crews had a special treat on the menu for dinner – one crew member cooked a special batch of Borscht for everyone. It was an incredible dish and a great opportunity to share our favorite dishes while on Mars. Following dinner, everyone came together for some down time before tucking in for the night at our respective habs.

Morgan Kainu, BA; MAU Journalist Crew 001

Sol Summary – January 29th

Crew 220 (MDRS) Sol Summary 29Jan2020

Sol: 10

Summary Title: Sweet Homeostasis

Author’s name: Connie Delisle

Mission Status: Nominal

Sol Activity Summary: The Crew rose at 08:00 to a cloudy sky with very light snow. Into mid-morning Mother Nature decided to bless us with high cloud and sun. MDRS again served as a relay point to MAU to communicate with the MAU Director. However, XO Wise has made good progress today in setting up a test WIFI connection which is part of the objectives set during this mission. The remaining Crew at MDRS, myself and astronomy researcher Marufa Akhter Bhuiyan thoroughly cleaned and organized the lower deck – floor, walls, airlocks, bathroom, cabinets. A safe, orderly homebase not only helps avoid accidents – it creates a sense of ownership and in becoming a home we could be forever living here on Mars if we didn’t have to return to Earth.

The MDRS Crew stayed close to home today to work in the GreenHab, planting radishes and romain lettuce. MDRS also served as HabComm to support the MAU Crew who participated in a 3-hour EVA focused on exploration and scouting. From a Raven’s eye view, the Team Leader remarked that “he felt part of the landscape of Mars”. The rolling red hills and blue sky was a remarkable backdrop to explore the soil, rocks and outcroppings.

Look Ahead Plan: Sol 11 includes an AM and PM EVA. These experiences aim to reinforce the medical training already received and introduce Search and Rescue (SAR). On Mars, medical emergencies may occur in the Hab or on EVA. The MAU-MDRS Crew testing and training is fundamental to establish a baseline understanding of the medical challenges in using “Earth” patient assessment protocols, conventional SAR equipment and applying operational protocols for recovery. All this hard work is for the purpose of innovating medical equipment, training and processes to improve the safety of security of astronauts on Mars.

Yesterday was a milestone in submitting reports before the CapComm window closure yesterday and to receive such positive comments on the photos. The Crew has done a great job on progressively becoming more organized, creating time and space to spend time on reports, science and to reflect on life on MARS. As this is my last night as Commander at MDRS, I am happy to transfer to Commander Hanacek, a station that I feel is well taken care of and appreciated.

Anomalies in work: None

Weather: Approximately -6C at night, max 3C during the day tomorrow; probability of precipitation is less than 20%.

Crew Physical Status: All the Crew is good health, bolstered after a good night’s sleep.

EVA: One morning and one in the afternoon.

Reports to be filed:

Journalist Report

Sol Summary

Photo Report
Operations Report
GreenHab Report

Support Requested: Revised request for two not four fluorescent bulbs for the lower deck. The engineer found two bulbs on the lower deck and has installed them.